OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of -1 and passes through the point (-2, 5).

OpenStudy (anonymous):

hey @jim_thompson5910 would you mind explaining to me how to solve these kinds of problems :)

jimthompson5910 (jim_thompson5910):

You would start with point-slope form, which is \[\Large y-y_{1} = m(x - x_{1})\] then plug in the given info to get \[\Large y-5 = -1(x - (-2))\] \[\Large y-5 = -1(x + 2)\]

OpenStudy (anonymous):

Thanks again @jim_thompson5910 you make geometry easy to understand